# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation in Low-Data Regimes

### v3

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth labels to compare and filter low-quality examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distill models in low-resource settings. To address this challenge, we propose a distillation framework that does not require any labeled data. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER compared to those without filtering and are on par with or perform better than similar supervised data filtering setups. When we scale the data, our models significantly outperform all zero-shot and supervised models. We demonstrate that it is possible to distill large Whisper models into relatively small ones without using any labeled data. Our distilled models are also 25-50% more compute- and memory-efficient while maintaining performance equal to or better than that of the teacher model.
@@ 1 Introduction
In an effort to democratize automatic speech recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023); Radford et al. (2023); Communication et al. (2023); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to high-resource counterparts such as English Radford et al. (2023). For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows the word error rate (WER) in the single digits on English and Spanish, a much lower performance is shown for East Asian and African languages [^1] https://github.com/openai/whisper/discussions/1762 Talafha et al. (2023). While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-intensive and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2020); Frantar et al. (2022).
@@
  • We assess the performance of our distilled models across diverse settings and extend the existing evaluation to include new datasets.
  • We analyze the effectiveness of our best metrics in detecting low-quality pseudo-labeled examples.
@@ 2 Related Work
To address the efficiency issue, knowledge distillation Hinton et al. (2015); Gou et al. (2021) is proven to be very effective. Different approaches for knowledge distillation have been explored such as Patient Knowledge Distillation (PKD) Sun et al. (2019), T-S learning Manohar et al. (2018), Cross-Modal Hashing (CMH) Hu et al. (2020), Joint Unsupervised Domain Adaptation with KD Zhang et al. (2021) and Data-Free KD Lopes et al. (2017) among others in the NLP and CV fields alike. Knowledge distillation has also been applied to speech recognition:  Shao et al. (2023) shrink a Whisper Radford et al. (2023) model tangibly ($\sim 80\%$ of its original size), all while improving its performance;  Chang et al. (2021) reduce a Hubert model by to 75% of its initial size through layer-wise distillation without a significant drop in performance.
Knowledge distillation has also been applied to speech-to-text tasks Nayem et al. (2023); Hentschel et al. (2024); Tian et al. (2022) to reduce the memory and compute requirements. More recently, it has been used to distill multilingual models from Whisper Ferraz et al. (2024) and strong monolingual models Gandhi et al. (2023), as well as in low-resource settings. An example of the latter is Waheed et al. (2024) who investigate pseudo-label distillation methods Gandhi et al. (2023) across various Modern Standard Arabic (MSA) and dialectal Arabic datasets. Their exploration demonstrate the efficacy of distillation in enhancing both efficiency and performance, showcasing that smaller dedicated models can outperform larger multilingual ones. While the pseudo-labeling approach has proven effective across diverse languages Gandhi et al. (2023); Waheed et al. (2024), it traditionally requires labeled data for filtering low-quality pseudo-labeled examples. In this work we address this limitation by introducing an unsupervised framework based on data filtering methods, thereby eliminating the dependency on labeled data altogether.
@@ 3.1 Knowledge Distillation
  • $\mathcal{L}_{PL}$ trains the student using pseudo labels as ground truth.
@@ 3.2 Label-Free Data Filtering
Previous data filtering methods, such as those used by Gandhi et al. (2023) and Waheed et al. (2024), rely on ground-truth labels by calculating WER between teacher-generated pseudo-labels and reference transcripts, discarding examples with high WER. However, this dependence on labeled data limits their use in low-resource languages where such labels are scarce.
@@ 3.2.1 Proxy Models
We use a trained ASR model, SeamlessM4T-large-v2 Communication et al. (2023), as a proxy to generate reference transcripts for the input speech. Then, the quality of pseudo-labels generated by the teacher model is evaluated by calculating the WER between the proxy’s and the teacher’s outputs, referred to as pWER. Lower pWER values indicate higher agreement and examples exceeding a defined pWER threshold are discarded.
3.3 Uncertainty Quantification
@@ 3.3.2 Multimodal Embeddings
3.3.3 Similarity with Synthetic Speech
Synthetic speech is generated from pseudo-labels using XTTS-v2222https://huggingface.co/coqui/XTTS-v2 and compared with input speech. Similarity is assessed via PESQ and MCD, where high PESQ and low MCD scores indicate higher label accuracy.
@@ 3.4 Training Data
We follow Waheed et al. (2024) for the data mixture to train our models. As a result, we randomly select segments from a diverse set of datasets, including MGB2 Ali et al. (2019a), MGB3 Ali et al. (2017), FLEURS Conneau et al. (2022), CommonVoice 15.0 Ardila et al. (2020), QASR Mubarak et al. (2021), the Arabic Speech Corpus Halabi (2016), and the Massive Arabic Speech Corpus (MASC) Al-Fetyani et al. (2021). Specifically, we sample 100K and 500K segments, equivalent to approximately 100 and 500 hours of pseudo-labeled speech data, respectively. Our dataset compilation strictly includes the train splits only from each source. We filter out roughly 27% bad quality examples from our training data using the metrics described in Section 3.
@@ 3.5 Teacher and Student Models
In line with previous work Gandhi et al. (2023); Waheed et al. (2024), we utilize the whisper-large-v2 checkpoint [^3] https://huggingface.co/openai/whisper-large-v2 for both pseudo-labeling and as the teacher model during training. We train two variants of students, differing in the number of layers removed from the teacher model. Following Gandhi et al. (2023) and Waheed et al. (2024), we initialize the student models with maximally spaced layers in the encoder and decoder block of the teacher model. We refer to the models with 16-16 and 32-16 encoder-decoder blocks as UDW-16-16 and UDW-32-16, respectively, when we train with 100K segments and as UDW-16-16++ and UDW-32-16++ when we train with 500K segments.
@@ 4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to include the SADA Alharbi et al. (2024) dataset. We describe our evaluation datasets below.
@@ 4.1 Evaluation Dataset
Multi-Genre Broadcast. The MGB dataset Ali et al. (2019a, 2017, b) consists of MGB2 (1,200 hours of MSA-dominant Aljazeera Arabic speech), MGB3 (6 hours of Egyptian dialectal speech), and MGB5 (14 hours of Moroccan dialectal speech).
@@ 4.2 Baselines
We evaluate several multilingual and monolingual speech recognition models on different Arabic varieties, including standard and accented MSA, as well as various dialects. The models are grouped as follows:
@@
We evaluate three baselines: Wav2Vec2-XLS-R Conneau et al. (2020); Babu et al. (2021) trained on CV8.0, HuBERT Hsu et al. (2021) trained on MGB-3 Ali et al. (2017) and a 5.5-hour Egyptian Arabic corpus, and a fine-tuned Whisper-large-v2 model trained on CV11.0 and MGB-2. These models have been used off the shelf and been trained by others.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we evaluate two distilled model variants of different sizes. We extend the evaluation to SADA’s test and valid splits, using 16-16 and 32-16 models trained with a WER threshold of 80% on both 100K and 500K segments. Additionally, we include models trained without filtering as a lower bound. For fair comparison with supervised data filtering, we follow an experimental setup identical to prior work Waheed et al. (2024).
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless specified otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and use WER and CER as the primary metrics for quantitative evaluation of the models.
For distillation, we select roughly 27% of the top segments based on the metrics, to match the amount of data used in a supervised setup when $\lambda=80\%$. We do not conduct hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023). The distillation process and its key parameters are detailed in Table 6 in Appendix C.
@@ 5 Results and Discussion
$\quad$-- DW-16-16 0.80 22.8/7.6 26.1/11.2 54.1/29.7 95.1/59.1 17.6/6.0 93.4/51.6 51.8/20.2 64.7/27.3 68.0/34.0 78.3/38.8 78.4/49.6
$\quad$-- DW-32-16 1.12 21.2/7.3 22.8/10.4 51.3/30.8 90.5/58.8 14.9/4.9 87.3/63.2 47.6/20.0 57.1/24.9 63.9/35.6 73.0/50.9 70.7/53.6
UDW-16-16 (0.80)
$\quad$-- nnll 23.59/8.15 26.43/11.26 54.37/27.98 92.25/55.25 18.49/6.26 82.44/41.4 63.33/25.7 52.97/20.52 72.03/35.96 93.4/49.2 84.3/55.0
$\quad$-- pesq 24.28/8.41 28.26/12.11 54.16/27.69 93.72/54.88 19.71/6.89 81.54/40.34 64.92/27.41 52.94/20.16 69.67/32.55 86.96/44.17 79.7/50.1
$\quad$-- entropy 23.84/8.1 27.41/12.17 57.29/31.24 94.75/56.64 18.16/6.4 90.05/48 54.76/22.81 64.26/27.67 73.72/37.85 96.35/52.56 92.0/61.8
$\quad$-- conf 23.47/7.83 27.2/11.87 52.66/27.85 87.79/50.73 17.66/6.12 89.15/43.94 53.12/20.29 63.44/25.52 69.03/31.75 79.11/39.27 78.6/49.2
$\quad$-- proxy 22.54/7.48 25.48/11.39 49.41/26.36 84.75/49.97 17.59/7.5 74.53/42.15 61.07/23.69 50.5/19.66 65.43/30.93 84.12/41.94 68.6/46.2
$\quad$-- sonar 24.06/8.04 28.34/11.86 55.09/28.66 85.59/49.21 20.42/7.06 86.93/43.48 56.36/22.61 65.4/27.43 70.7/32.89 76.48/36.13 74.9/45.6
@@
$\quad$-- nnll 18.92/6.24 22.63/10.12 48.08/25.39 94.07/55.53 13.29/4.47 75.43/35.85 55.24/20.88 44.69/16.04 63.71/30.49 83.81/41.38 73.4/46.1
$\quad$-- pesq 21.7/7.5 24.2/10.6 49.0/26.4 87.0/51.0 16.1/5.3 88.8/43.2 47.5/17.2 58.5/22.7 67.0/30.9 75.5/36.4 73.3/48.1
$\quad$-- entropy 19.09/6.53 21.82/10.34 46.93/28.71 87.11/66.87 13.29/4.34 90.99/84.02 46.68/21.07 58.16/31.08 64.94/44.23 76.35/54.25 68.3/52.3
$\quad$-- conf 20.27/6.46 22.76/9.79 46.6/23.42 83.43/48.61 14.53/4.73 73.42/36.08 57.25/21.99 61.06/25.82 47.5/17.7 85.71/42.82 66.2/41.3
$\quad$-- proxy 19.04/6.17 21.51/9.87 44.29/22.45 82.64/46.05 14.23/6.23 80.36/36.11 45.46/15.6 56.01/20.88 61.63/25.69 69.52/29.49 64.6/41.9
$\quad$-- sonar 17.82/5.62 21.1/8.98 45.29/23.4 80.39/46.97 13.07/4.41 78.97/37 44.3/15.11 54.34/19.56 58.76/24.29 66.82/28.24 58.7/35.6
@@
$\quad$-- proxy 15.5/4.8 23.3/10.4 46.8/24.3 84.6/48.0 14.2/4.8 82.9/41.6 46.4/16.3 57.0/21.1 61.1/27.5 70.8/33.0 60.2/35.3
$\quad$-- sonar 18.9/6.1 22.8/9.8 47.2/24.2 82.0/47.1 14.9/5.0 81.5/38.9 47.7/17.2 57.4/21.4 62.1/26.2 69.2/29.5 59.4/33.96
UDW-32-16-++ (1.12)
$\quad$-- proxy 18.1/5.8 20.8/9.2 43.1/21.1 82.3/44.1 12.6/4.2 80.8/37.1 43.1/14.5 54.4/20.1 57.7/23.1 68.5/29.2 55.1/31.4
$\quad$-- sonar 17.0/5.3 21.9/9.9 44.1/22.4 79.8/44.2 12.8/3.9 78.2/34.6 54.2/19.2 44.4/15.1 58.2/23.2 66.6/27.0 61.6/33.4
@@
Table 1: WER/CER scores after normalization and removing diacritics. All baseline distilled models (dw-) are trained with a filtering threshold of 80 if not specified. Best results are shown in bold. Second best results are underlined. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, U - Unsupervised, D - Distil.
@@
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. We distil two variants of student models. The first student has 16 of each of the encoder and decoder blocks (UDW-16-16) and the second one has 32 encoder and 16 decoder blocks (UDW-32-16). We experiment with 100K and 500K segments. We report our main results in Table 1, and the average across different evaluation setups in Table 2. Results on the top five varieties [^4] We refer labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 3. The reported results are computed on normalized text that includes removing diacritics. Results on the text with no normalization can be found in Appendix 3.2.
@@
In zero-shot models, we compare our results with MMS Pratap et al. (2023), three variants of SeamlessM4T family models (medium, large-v1, large-v2), and four Whisper models. We extend the evaluation to the SADA data for these models as well. While larger variants of both SeamlessM4T and Whisper perform well on standard benchmark setups like Common Voice and Fleurs, they perform poorly on novel and challenging settings, such as dialectal speech. For example, the best zero-shot model (SeamlessM4T-large-v2) has a 7.6% WER on Fleurs compared to 65.0% on the SADA test set as shown in Table 1. This underscores the inadequate evaluation and limited generalization capability of zero-shot models in unseen and challenging settings.
As demonstrated in the reported results, the best-distilled model DW-32-16++ outperforms all other models including its teacher (large-v2), the best Whisper (large-v3), and the best baseline (SeamlessM4T-large-v2).
We train two versions of teacher models using datasets of size 100K and 500K segments, respectively. Approximately 27% of the examples were removed from each dataset based on specific filtering metrics, matching the amount of data used in the supervised setup when the filtering threshold is 80.
@@
From our experiments, we observe that two of our best filtering measures, sonar-sim and proxy-ref, perform similar to or surpass the equivalent supervised setups. For instance, DW-32-16 trained on 100K examples with a filtering threshold of 80 achieves a WER of 35.3% on the benchmark test split, 66.6% on the SADA test split, and 61.1% averaged on five novel dialects. By comparison, our best-performing model in a similar configuration, UDW-32-16-sonar, achieves 35.5% WER on the benchmark test split, 58.7% on the SADA test, and 60.6% on five dialects. This reflects close to 3% improvement in average score compared to the supervised setup (51.6 vs 54.3) and more than a 7% improvement over the no-filter setup, as detailed in Table 2.
In addition to that, our experiments show that the smaller student model (UDW-16-16) performs better with proxy-ref as a filtering measure. More specifically, the UDW-16-16, experimented with 100K examples and utilizing proxy-ref to filter out low-quality pseudo labels, and matches the results of supervised setups (58.5% vs 57.6%) and shows over 6% improvement compared to setups where pseudo-labels are not filtered (64.2%).
@@
$\quad$-- proxy 40.0/20.5 42.3/22.1 68.6/46.2 65.4/42.2 67.1/31.7
$\quad$-- sonar 42.7/21.0 44.9/22.8 74.9/45.6 71.2/43.5 71.2/32.5
@@
$\quad$-- proxy 36.3/18.2 39.1/19.6 64.6/41.9 58.7/36.1 62.6/25.6
$\quad$-- sonar 35.5/17.9 37.8/19.7 58.7/35.6 56.4/34.7 60.6/24.8
@@
$\quad$-- proxy 36.9/18.4 41.4/21.2 64.1/39.2 60.2/35.3 63.6/27.9
$\quad$-- sonar 37.2/18.4 39.3/19.9 61.7/35.4 59.4/34.0 63.6/26.6

### v4

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation using pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth labels to compare with and filter low-quality examples, making the process dependent on human labels. Additionally, the distillation process requires a large amount of data thereby limiting its applicability in low-resource settings. To address this, we propose a distillation framework that does not require any labeled data. Through experimentation, we show that our best-distilled models outperform the teacher model by 5-7 WER points and are on par with or outperform similar supervised data filtering setups. When scaling the data, our models significantly outperform all zero-shot and supervised models. Our models are also 25-50% more compute- and memory-efficient while maintaining performance equal to or better than that of the teacher model. For more details about our models, dataset, and other resources, please visit our GitHub page: https://github.com/UBC-NLP/uDistilWhisper.
@@ 1 Introduction
To democratize automatic speech recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023); Radford et al. (2023); Communication et al. (2023); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to high-resource counterparts such as English Radford et al. (2023). For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows the word error rate (WER) in the single digits on English and Spanish, a much lower performance is shown for East Asian and African languages [^1] https://github.com/openai/whisper/discussions/1762 Talafha et al. (2023, 2024). While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020). In addition to the disparity in performance, multilingual models are extremely compute-intensive and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2020); Frantar et al. (2022).
@@
  • We assess the performance of our distilled models, which effectively outperform the teacher model by 5–7 WER points, and extend the existing setup to include new datasets in both Arabic and Swahili.
  • We analyze the effectiveness of our best metrics in detecting low-quality pseudo-labeled examples, achieving an AUC ranging between $0.77$ and $0.82$ for examples with WER > 20.
@@ 2 Related Work
To address the efficiency issue, knowledge distillation Hinton et al. (2015); Gou et al. (2021) is proven to be very effective. Different approaches for knowledge distillation have been explored such as Patient Knowledge Distillation (PKD) Sun et al. (2019), T-S learning Manohar et al. (2018), Cross-Modal Hashing (CMH) Hu et al. (2020), Joint Unsupervised Domain Adaptation with KD Zhang et al. (2021) and Data-Free KD Lopes et al. (2017) among others in the NLP and CV fields alike. Knowledge distillation has also been applied to speech recognition:  Shao et al. (2023) shrink a Whisper Radford et al. (2023) model tangibly ($\sim 80\%$ of its original size), all while improving its performance;  Chang et al. (2022) reduce a Hubert model to 75% of its initial size through layer-wise distillation without a significant drop in performance.
In addition, knowledge distillation has found successful applications in speech-to-text tasks as well,  Nayem et al. (2023); Hentschel et al. (2024); Tian et al. (2022) effectively reducing the memory and compute requirements. More recently, it has been used to distill multilingual models from Whisper Ferraz et al. (2024) and strong monolingual models Gandhi et al. (2023), as well as in low-resource settings. An example of the latter is Waheed et al. (2024) who investigate pseudo-label distillation methods Gandhi et al. (2023) across various Modern Standard Arabic (MSA) and dialectal Arabic datasets. Their exploration demonstrate the efficacy of distillation in enhancing both efficiency and performance, showcasing that smaller dedicated models can outperform larger multilingual ones. While the pseudo-labeling approach has proven effective across diverse languages Gandhi et al. (2023); Waheed et al. (2024), it traditionally requires labeled data for filtering low-quality pseudo-labeled examples. In this work we address this limitation by introducing an unsupervised framework based on data filtering methods, thereby eliminating the dependency on labeled data altogether.
@@ 3.1 Knowledge Distillation
  • $\mathcal{L}_{PL}$ trains the student using pseudo-labels as ground truth.
@@ 3.2 Label-Free Data Filtering
Previous data filtering methods, such as those used by Gandhi et al. (2023) and Waheed et al. (2024), rely on ground-truth labels by computing the WER between teacher-generated pseudo-labels and reference transcripts, discarding examples with high error rates. However, this dependence on labeled data limits their use in low-resource languages where such labels are scarce.
@@ 3.2.1 Proxy Models
We use a pre-trained ASR model, SeamlessM4T-large-v2 Communication et al. (2023), as a proxy to generate reference transcripts for the input speech. Then, the quality of pseudo-labels generated by the teacher model is evaluated by calculating the WER between the proxy’s and the teacher’s outputs, referred to as pWER. Lower pWER values indicate higher agreement and examples exceeding a defined pWER threshold are discarded.
3.2.2 Uncertainty Quantification
@@ 3.2.4 Multimodal Embeddings
3.2.5 Perceptual Evaluation of Speech Quality (PESQ)
Synthetic speech is generated from pseudo-labels using XTTS-v2222https://huggingface.co/coqui/XTTS-v2 and compared with input speech. Similarity is assessed via PESQ, where high scores indicate higher label accuracy.
@@ 3.3 Training Data
We follow Waheed et al. (2024) for the data mixture to train our models. As a result, we randomly select segments from a diverse set of datasets, including MGB2 Ali et al. (2016), MGB3 Ali et al. (2017), FLEURS Conneau et al. (2023), CommonVoice 15.0 Ardila et al. (2020), QASR Mubarak et al. (2021), the Arabic Speech Corpus Halabi (2016), and the Massive Arabic Speech Corpus (MASC) Al-Fetyani et al. (2023). Specifically, we sample 100K and 500K segments, equivalent to approximately 100 and 500 hours of pseudo-labeled speech data, respectively. Our dataset compilation strictly includes the train splits only from each source. We filter out roughly 27% low quality examples from our training data using the metrics described in Section 3.
@@ 3.4 Teacher and Student Models
In line with previous work Gandhi et al. (2023); Waheed et al. (2024), we utilize the whisper-large-v2 checkpoint [^3] https://huggingface.co/openai/whisper-large-v2 for both pseudo-labeling and as the teacher model during training. We train two variants of students, differing in the number of layers removed from the teacher model. Following Gandhi et al. (2023) and Waheed et al. (2024), we initialize the student models with maximally spaced layers in the encoder and decoder block of the teacher model. When training on 100K segments, we refer to the models with 16-16 and 32-16 encoder-decoder blocks as UDW-16-16444Following the format from Waheed et al. (2024), UDW refers to Unsupervised Distill Whisper. and UDW-32-16, respectively. Similarly, we refer to the models trained on 500K segments as UDW-16-16++ and UDW-32-16++.
@@ 4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, we extend the evaluation to include the SADA Alharbi et al. (2024) and Casablanca [^5] The In-House data in Waheed et al. (2024), reported in Table 5, is an earlier subset of the Casablanca dataset, used prior to the official release of the larger version. Talafha et al. (2024) datasets. We describe all the used datasets below.
@@ 4.1 Evaluation Dataset
Multi-Genre Broadcast. The MGB dataset Ali et al. (2016, 2017, 2019) consists of MGB2 (1,200 hours of MSA-dominant Aljazeera Arabic speech), MGB3 (six hours of Egyptian dialectal speech), and MGB5 (14 hours of Moroccan dialectal speech).
@@
Casablanca. Casablanca Talafha et al. (2024) is a dialectal Arabic dataset that contains $48$ hours of speech from various TV shows. It covers eight different dialects and is annotated by native speakers. The results on this dataset are separately reported in Table 5.
@@ 4.2 Baselines
We employ several monolingual and multilingual speech recognition models on different Arabic varieties, including standard and accented MSA, as well as various dialects. The models are grouped as follows:
@@
We evaluate three baselines, namely Wav2Vec2-XLS-R Conneau et al. (2020); Babu et al. (2022) trained on CV8.0, HuBERT Hsu et al. (2021) trained on both MGB-3 Ali et al. (2017) and a 5.5-hour Egyptian Arabic corpus, and a fine-tuned Whisper-large-v2 model trained on CV11.0 and MGB-2. We use these models off-the-shelf from their respective checkpoints.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, two distilled model variants of different sizes are evaluated. We extend the evaluation to SADA’s test and valid splits, using 16-16 and 32-16 models trained with a WER threshold of 80% on both 100K and 500K segments. Additionally, we include models trained without filtering as a lower bound. For a fair comparison with supervised data filtering, we follow an experimental setup identical to prior work Waheed et al. (2024).
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless specified otherwise. Throughout our experiments, a maximum sequence length of 225 is maintained and WER and CER are used as the primary metrics for the quantitative evaluation of the models.
For distillation, roughly 27% of the top segments are selected based on the metrics, to match the amount of data used in a supervised setup when $\lambda=80\%$. We do not conduct hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023). The distillation process and its key parameters are detailed in Table 13 in Appendix C.
@@ 5 Results and Discussion
$\quad$-- DW-16-16 0.80 22.8 26.1 54.1 95.1 17.6 93.4 51.8 64.7 68.0 78.3 78.4
$\quad$-- DW-32-16 1.12 21.2 22.8 51.3 90.5 14.9 87.3 47.6 57.1 63.9 73.0 70.7
Ours
UDW-16-16 0.80
$\quad$-- nll 23.6 26.4 54.4 92.3 18.5 82.4 63.3 53.0 72.0 93.4 84.3
$\quad$-- pesq 24.3 28.3 54.2 93.7 19.7 81.5 64.9 52.9 69.7 87.0 79.7
$\quad$-- entropy 23.8 27.4 57.3 94.8 18.2 90.1 54.8 64.3 73.7 96.4 92.0
$\quad$-- conf 23.5 27.2 52.7 87.8 17.7 89.2 53.1 63.4 69.0 79.1 78.6
$\quad$-- proxy 22.5 25.5 49.4 84.8 17.6 74.5 61.1 50.5 65.4 84.1 68.6
$\quad$-- sonar 24.1 28.3 55.1 85.6 20.4 86.9 56.4 65.4 70.7 76.5 74.9
@@
$\quad$-- nll 18.9 22.6 48.1 94.1 13.3 75.4 55.2 44.7 63.7 83.8 73.4
$\quad$-- pesq 21.7 24.2 49.0 87.0 16.1 88.8 47.5 58.5 67.0 75.5 73.3
$\quad$-- entropy 19.1 21.8 46.9 87.1 13.3 91.0 46.7 58.2 64.9 76.4 68.3
$\quad$-- conf 20.3 22.8 46.6 83.4 14.5 73.4 57.3 61.1 47.5 85.7 66.2
$\quad$-- proxy 19.0 21.5 44.3 82.6 14.2 80.4 45.5 56.0 61.6 69.5 64.6
$\quad$-- sonar 17.8 21.1 45.3 80.4 13.1 79.0 44.3 54.3 58.8 66.8 58.7
@@
$\quad$-- proxy 15.5 23.3 46.8 84.6 14.2 82.9 46.4 57.0 61.1 70.8 60.2
$\quad$-- sonar 18.9 22.8 47.2 82.0 14.9 81.5 47.7 57.4 62.1 69.2 59.4
UDW-32-16-++ 1.12
$\quad$-- proxy 18.1 20.8 43.1 82.3 12.6 80.8 43.1 54.4 57.7 68.5 55.1
$\quad$-- sonar 17.0 21.9 44.1 79.8 12.8 78.2 54.2 44.4 58.2 66.6 61.6
@@
Table 1: WER ($\downarrow$) after normalization and removing diacritics. All baseline distilled models (dw-) are trained with a filtering threshold of 80 if not specified. Best results are shown in bold. Second best results are underlined. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, DW - Distill Whisper, UDW - Unsupervised Distill Whisper, nll - negative log likelihood, conf - confidence score.
@@
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. Our main results can be found in Table 5, and the average across different evaluation setups in Table 5. Results on the top five varieties [^6] We refer labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 3.1. The reported results are computed on normalized text that includes removing diacritics. Results on the text with no normalization can be found in Appendix 3.3.
@@
While larger variants of both SeamlessM4T and Whisper perform well on standard benchmark setups like Common Voice and Fleurs, we find that they perform poorly on dialectal speech. For example, the best zero-shot model (SeamlessM4T-large-v2) has a 7.6% WER on Fleurs compared to 65.0% on the SADA test set as shown in Table 5. This underscores the inadequate evaluation and limited generalization capability of zero-shot models in unseen and challenging settings. The best-distilled model DW-32-16++ outperforms all other models including its teacher (large-v2), the best Whisper version (large-v3), and the best baseline (SeamlessM4T-large-v2).
@@
When it comes to our distilled models, we observe that the two best filtering measures, sonar-sim and proxy-ref, are comparable to or surpass the equivalent supervised setups. For instance, DW-32-16 trained on 100K examples with a filtering threshold of 80 achieves a WER of 35.3% on the benchmark test split, 66.6% on the SADA test split, and 61.1% averaged on five novel dialects. By comparison, our best-performing model in a similar configuration, UDW-32-16-sonar, achieves 35.5% WER on the benchmark test split, 58.7% on the SADA test, and 60.6% on five dialects. This reflects close to 3% improvement in average score compared to the supervised setup (51.6 vs 54.3) and more than a 7% improvement over the no-filter setup, as detailed in Table 5. In addition to that, our experiments show that the smaller student model (UDW-16-16) performs better with the proxy-ref method. More specifically, when trained on 100K examples, it matches the results of supervised setups (58.5% vs 57.6%) and shows over 6% improvement compared to setups where pseudo-labels are not filtered (64.2%).
@@
$\quad$-- proxy 40.0 42.3 68.6 65.4 67.1
$\quad$-- sonar 42.7 44.9 74.9 71.2 71.2
@@
$\quad$-- proxy 36.3 39.1 64.6 58.7 62.6
$\quad$-- sonar 35.5 37.8 58.7 56.4 60.6
@@
$\quad$-- proxy 36.9 41.4 64.1 60.2 63.6
$\quad$-- sonar 37.2 39.3 61.7 59.4 63.6